# OmniHD-Scenes: A Next-Generation Multimodal Dataset for Autonomous Driving

### v4

@@ Abstract
The rapid advancement of deep learning has intensified the need for comprehensive data for use by autonomous driving algorithms. High-quality datasets are crucial for the development of effective data-driven autonomous driving solutions. Next-generation autonomous driving datasets must be multimodal, incorporating data from advanced sensors that feature extensive data coverage, detailed annotations, and diverse scene representation. To address this need, we present OmniHD-Scenes, a large-scale multimodal dataset that provides comprehensive omnidirectional high-definition data. The OmniHD-Scenes dataset combines data from 128-beam LiDAR, six cameras, and six 4D imaging radar systems to achieve full environmental perception. The dataset comprises 1501 clips, each approximately 30-s long, totaling more than 450K synchronized frames and more than 5.85 million synchronized sensor data points. We also propose a novel 4D annotation pipeline. To date, we have annotated 200 clips with more than 514K precise 3D bounding boxes. These clips also include semantic segmentation annotations for static scene elements. Additionally, we introduce a novel automated pipeline for generation of the dense occupancy ground truth, which effectively leverages information from non-key frames. Alongside the proposed dataset, we establish comprehensive evaluation metrics, baseline models, and benchmarks for 3D detection and semantic occupancy prediction. These benchmarks utilize surround-view cameras and 4D imaging radar to explore cost-effective sensor solutions for autonomous driving applications. Extensive experiments demonstrate the effectiveness of our low-cost sensor configuration and its robustness under adverse conditions. Data will be released at https://www.2077ai.com/OmniHD-Scenes.
@@ 1 Introduction
TABLE I: Comparison with publicly datasets for autonomous driving, where “Scenes” represents the number of consecutive clips and “/” represents unsupported or no statistics. “Weather” represents adverse conditions such as rain or fog. “d/n” denotes day and night, and “CTS” denotes Closed Test Site data. Bolded line represents OmniHD-Scenes.
Dataset Year Anno. Scenes Sensors 3D Box Tasks Whether Time CTS
@@
MAN TruckScenes[20] 2024 360° 747 4x 6x4d 6×64-beam 1.0M $\checkmark$ $\checkmark$ $\times$ $\checkmark$ d/n $\checkmark$
@@ 2.1 Autonomous Driving Dataset
While these vision-based methods yield promising results, LiDAR point clouds provide precise geometric and depth information that can address the inherent limitations of camera-based 3D detection. As a result, some methods have achieved excellent 3D detection performance using LiDAR point clouds [28, 29, 30]. Others have achieved enhanced multimodal detection performance through the use of fusion approaches [31, 32]. In particular, BEVFusion [31] has an effective framework that integrates LiDAR and multi-view camera data. This framework transforms camera features to BEV space using an LSS-based approach that is independent of LiDAR processing. Hence, robust multimodal detection is enabled through flexible sensor fusion.
Additionally, radar sensors offer a cost-effective alternative and have unique advantages in adverse weather conditions [33]. In the field of camera and radar fusion-based object detection, some methods focus on ROI-based approaches [34, 35, 36, 37, 38]. However, recent mainstream techniques have increasingly adopted BEV representations [39, 40, 41]. Furthermore, with the emergence of 4D radar, research has increasingly focused on the leveraging of 4D-radar features for detection tasks. Some recent methods utilize 4D-radar tensors as input to preserve more raw signal information [17]; however, this approach increases the computational resource consumption. Most studies have concentrated on the leveraging of 4D-radar point clouds as input for detection tasks. For example, in the RPFA-Net [42] framework, the radar point feature attention module was introduced to extract global contextual features from radar point clouds. In the SMURF [43] framework, the sparsity and noise of 4D-radar point clouds are mitigated through the use of pillarization and density features derived from the kernel density estimation of multi-dimensional Gaussian mixtures. RCFusion [44] leverages RadarPillarNet for hierarchical feature extraction from 4D-radar point clouds, and effectively fuses the BEV features of camera and 4D-radar modalities using interactive attention. Finally, LXL [45] leverages radar BEV features to infer a 3D occupancy network and assists image-based view transformation using 3D occupancy and estimated depth.
@@ 3.1 Sensor Suite
LiDAR Robosense RS-Ruby-128 ($\times$1) 128 beams, $360^{\circ}$ HFOV, $-25^{\circ}\sim+15^{\circ}$ VFOV, 10 FPS
Camera SG8-AR0820C ($\times$1) Front-view, $60^{\circ}$ HFOV, $33^{\circ}$ VFOV, 3840$\times$2160 resolution, 30 FPS
SG8-AR0820C ($\times$1) Rear-view, $120^{\circ}$ HFOV, $55^{\circ}$ VFOV, 3840$\times$2160 resolution, 30 FPS
SG2-AR0233C ($\times$4) Surround-view, $100^{\circ}$ HFOV, $53^{\circ}$ VFOV, 1920$\times$1080 resolution, 30 FPS
4D Radar GPAL Ares-R7861 ($\times$2) Front/Rear-view, $76\sim 79$ GHz, $140^{\circ}$/$20^{\circ}$ HFOV, $24^{\circ}$/$10^{\circ}$ VFOV, 13 FPS
@@
Domain Controller TZTEK-GEACX2 ($\times$1) 2$\times$ Jetson AGX Orin Processors, 400TOPS
@@ 3.5.3 Scene Selection and Anonymization
2:$\mathcal{K}\leftarrow\text{SelectKeyFrames}(\mathcal{S})$
3:$\mathcal{B}_{init}\leftarrow\mathcal{L}_{pre}(\mathcal{K})$
4:$\mathcal{B}_{gt}\leftarrow\text{ManualRefineBBoxwithID}(\mathcal{K},\mathcal{B}_{init})$
5:$\mathcal{A}_{final}\leftarrow\emptyset$
@@
16:         if $conf>max\_conf$ then
@@
34:$\mathcal{A}_{final}\leftarrow\text{Check}(\mathcal{A}_{final})$
@@ 3.6.2 Static Point Cloud Segmentation Annotation
4:$P_{r}\leftarrow\text{RoadSeg}(P_{\text{st}},M)$
5:$P_{n}\leftarrow P_{\text{st}}\setminus P_{r}$
@@
12:$S_{\text{global}}\leftarrow\text{TransformToGlobal}(S_{\text{static}},\mathcal{P})$
13:$S_{\text{labeled}}\leftarrow\text{UseSegLabel}(S_{\text{global}})$
@@
24:$S_{\text{local}}\leftarrow\text{TransformToLocal}(S_{\text{labeled}},\mathcal{P}^{-1})$
25:$S_{\text{combined}}\leftarrow\text{CombineScenes}(D_{\text{agg}},S_{\text{local}})$
26:$S_{\text{filtered}}\leftarrow\text{RadiusFilter}(S_{\text{combined}})$
@@
28:$\mathcal{O}\leftarrow\text{NN}(\text{Voxelization}(S_{\text{filtered}}))$
@@ 4 Benchmark Experiments
We established robust benchmarks for 3D object detection and 3D occupancy prediction tasks using the OmniHD-Scenes dataset. Following standard practice, we partitioned the annotated sequences into training and test sets at a ratio of 7:3 to facilitate model development and benchmark evaluation. We implemented several state-of-the-art baseline models, providing thorough evaluation metrics and performing in-depth result analysis. Our aim was to establish strong baseline performance standards that would facilitate and advance future research efforts.
@@ 4.1.2 Baselines and Implementation Details
As the 4D-radar baseline, we chose pillar-based methods; namely, PointPillars [28] and RadarPillarNet [44]. The input point-cloud range was {$(x,y,z)$||||$-60m<x<60m$, $-40m<y<40m$, $-3m<z<5m$}. The input raw radar point-cloud features $f_{in}$ were as follows:
@@
where $x,y,z$ are the coordinates; $v_{xr}$ and $v_{yr}$ are the radial velocity components compensated for ego-motion; $Power$ and $SNR$ are the amplitude and signal-to-noise ratio of the radar point cloud, respectively; and $t_{diff}$ is the time difference from the first frame. By default, we used three frames of accumulated radar points as input. The pillar size for both was $[0.25,0.25,-8]m$, and both were trained using the AdamW optimizer with an initial learning rate of 1e-3 for 24 epochs.
@@ 4.2.1 Evaluation Metrics
where $TP$, $FP$, and $FN$ represent the true-positive, false-positive, and false-negative prediction counts, respectively. Further, $C$ denotes the number of categories, excluding free space. Additionally, we assessed the occupancy-state estimation accuracy using the scene completion IoU (SC IoU), which evaluates geometric accuracy by computing IoU metrics for free space and occupied elements.
@@ 5 Conclusion and Future Work
In this study, we introduced OmniHD-Scenes, a large-scale multimodal dataset for exploration of the application of multi-view cameras and 4D radar as cost-effective sensors for autonomous driving. The dataset comprises 1,501 continuous 30-s clips and more than 450K synchronized frames, which capture complex traffic scenarios across various weather conditions and times, including extreme test cases from closed test sites. We presented an advanced 4D annotation pipeline that provides detailed annotations for 200 clips, including more than 514K 3D tracking boxes and static-scene point-cloud segmentation annotations. Additionally, we introduced a novel 3D semantic occupancy generation pipeline that automatically creates dense occupancy labels by leveraging the rich annotation data obtained previously. Using OmniHD-Scenes, we established benchmarks for 3D object detection and 3D semantic occupancy prediction. Comprehensive evaluation of mainstream algorithms elucidated their perception performance under adverse conditions. The experimental results demonstrated the potential of 4D-radar applications and are expected to motivate the development of higher-efficiency algorithms.
In the future, we plan to expand OmniHD-Scenes to facilitate additional tasks and benchmarks, including trajectory prediction, visual grounding, and end-to-end autonomous driving.
@@ Acknowledgement
The first authors thank to Intelligent Connected Vehicle inspection Center (Hunan) of CAERI Co., Ltd and Shanghai Geometrical Perception and Learning Co., Ltd.

### v5

@@ Abstract
The rapid advancement of deep learning has intensified the need for comprehensive data for use by autonomous driving algorithms. High-quality datasets are crucial for the development of effective data-driven autonomous driving solutions. Next-generation autonomous driving datasets must be multimodal, incorporating data from advanced sensors that feature extensive data coverage, detailed annotations, and diverse scene representation. To address this need, we present OmniHD-Scenes, a large-scale multimodal dataset that provides comprehensive omnidirectional high-definition data. The OmniHD-Scenes dataset combines data from 128-beam LiDAR, six cameras, and six 4D imaging radar systems to achieve full environmental perception. The dataset comprises 1501 clips, each approximately 30-s long, totaling more than 450K synchronized frames and more than 5.85 million synchronized sensor data points. We also propose a novel 4D annotation pipeline. To date, we have annotated 200 clips with more than 514K precise 3D bounding boxes. These clips also include semantic segmentation annotations for static scene elements. Additionally, we introduce a novel automated pipeline for generation of the dense occupancy ground truth, which effectively leverages information from non-key frames. Alongside the proposed dataset, we establish comprehensive evaluation metrics, baseline models, and benchmarks for 3D detection and semantic occupancy prediction. These benchmarks utilize surround-view cameras and 4D imaging radar to explore cost-effective sensor solutions for autonomous driving applications. Extensive experiments demonstrate the effectiveness of our low-cost sensor configuration and its robustness under adverse conditions. The dataset is available at https://www.2077ai.com/OmniHD-Scenes.
@@ 1 Introduction
TABLE I: Comparison with public datasets for autonomous driving, where “Scenes” represents the number of consecutive clips and “/” represents unsupported or no statistics. “Weather” represents adverse conditions such as rain or fog. “d/n” denotes day and night, and “CTS” denotes Closed Test Site data. Bolded line represents OmniHD-Scenes.
Dataset Year Anno. Scenes Sensors 3D Box Tasks Weather Time CTS
@@
MAN TruckScenes [8] 2024 360° 747 4x 6x4d 6×64-beam / $\checkmark$ $\checkmark$ $\times$ $\checkmark$ d/n
@@ 2.1 Autonomous Driving Dataset
While vision-based methods show promise, LiDAR provides precise geometric information that addresses camera inherent limitations in 3D detection. Several methods have achieved excellent performance using either LiDAR point clouds [17, 55, 39] or multimodal fusion approaches [23, 28]. Notably, BEVFusion [23] effectively integrates LiDAR and camera data through LSS-based camera-to-BEV transformation, enabling robust multimodal detection through flexible sensor fusion.
Additionally, radar sensors offer cost-effective advantages, particularly in adverse weather conditions [54]. While early camera-radar fusion methods focused on ROI-based approaches [32, 46], recent techniques have shifted toward BEV representations [16, 50]. With the emergence of 4D radar, studies have explored point cloud-based approaches for 4D radar detection [52, 26]. Additionally, RCFusion [60] leverages RadarPillarNet for hierarchical feature extraction from 4D-radar point clouds, and effectively fuses the BEV features of camera and 4D-radar modalities using interactive attention. LXL [51] leverages radar BEV features to infer a 3D occupancy network and assists image-based view transformation using 3D occupancy and estimated depth.
@@ 3.1 Sensor Suite
LiDAR Robosense RS-Ruby-128 ($\times$1) 128 beams, $360^{\circ}$ HFOV, $25^{\circ}\sim+15^{\circ}$ VFOV, 10 FPS
[…]
4D Radar GPAL Ares-R7861 ($\times$2) Front/Rear-view, $76\sim 79$ GHz, $140^{\circ}$$20^{\circ}$ HFOV, $24^{\circ}$$10^{\circ}$ VFOV, 13 FPS
@@
[…]
@@ 3.5.3 Scene Selection and Anonymization
[…]
@@
16:   if conf>max_confconf>max\_conf then
@@
[…]
@@ 3.6.2 Static Point Cloud Segmentation Annotation
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 Benchmark Experiments
We established robust benchmarks for 3D object detection and 3D occupancy prediction tasks using the OmniHD-Scenes dataset. To facilitate standard benchmarking, the 200 annotated clips are partitioned into training, validation, and test sets at a 6:1:3 ratio. Considering the evaluation inaccuracies caused by scene overlap in online mapping [24], we ensured that 93.33% of our splits have no geographic overlap. For the downstream occupancy prediction task, the negligible remaining overlap does not negatively impact the evaluation, as the distinct dynamic objects and varying environmental conditions (e.g., weather and lighting) prevent memorization, thereby establishing a rigorous standard Our aim was to establish strong baseline performance standards that would facilitate and advance future research efforts.
@@ 4.1.2 Baselines and Implementation Details
As the 4D-radar baseline, we chose pillar-based methods; namely, PointPillars [17] and RadarPillarNet [60]. The input point-cloud range was {(x,y,z)(x,y,z)||60m<x<60m-60m<x<60m, 40m<y<40m-40m<y<40m, 3m<z<5m-3m<z<5m}. The input raw radar point-cloud features $f_{in}$ were as follows:
@@
where $x,y,z$ are the coordinates; $v_{xr}$ and $v_{yr}$ are the radial velocity components compensated for ego-motion; $Power$ and $SNR$ are the amplitude and signal-to-noise ratio of the radar point cloud, respectively; and $t_{diff}$ is the time difference from the first frame. By default, we used three frames of accumulated radar points as input. The pillar size for both was [0.25,0.25,8]m[0.25,0.25,-8]m, and both were trained using the AdamW optimizer with an initial learning rate of 1e-3 for 24 epochs.
@@ 4.2.1 Evaluation Metrics
where $TP$, FPFP, and $FN$ represent the true-positive, false-positive, and false-negative prediction counts, respectively. Further, $C$ denotes the number of categories, excluding free space. Additionally, we assessed the occupancy-state estimation accuracy using the scene completion IoU (SC IoU), which evaluates geometric accuracy by computing IoU metrics for free space and occupied elements.
@@ 5 Conclusion and Future Work
In this study, we introduced OmniHD-Scenes, a large-scale, multimodal dataset featuring multi-view cameras and 4D radar to explore cost-effective sensor solutions for autonomous driving. The dataset contains 1,501 clips of complex driving scenarios, including hazardous test cases from closed sites. Using an advanced 4D annotation pipeline, we annotated 200 of these clips with over 514,000 3D tracking boxes and static-scene point-cloud segmentation. Additionally, we introduced a novel 3D semantic occupancy generation pipeline that automatically creates dense occupancy labels. We established benchmarks for 3D object detection and occupancy prediction, with experiments demonstrating the potential of 4D radar in adverse conditions and aiming to inspire more efficient algorithm development. In the future, we plan to expand the OmniHD-Scenes dataset to support the development and evaluation of next-generation autonomous driving algorithms. Our efforts will be concentrated on the following areas.
End-to-End AEB Benchmark. We plan to establish an end-to-end AEB benchmark focused on safety-critical scenarios. The OmniHD-Scenes dataset includes hazardous test cases from closed test sites, featuring Autonomous Emergency Braking (AEB) scenarios triggered by specific Time-to-Collision (TTC) conditions, which makes an end-to-end AEB benchmark possible. An end-to-end AEB model can make decisions (e.g., braking) directly from raw sensor input, and its performance can be evaluated in controlled, repeatable, safety-critical contexts. Furthermore, this allows for a re-evaluation of the role of 4D radar in the system. The direct and precise velocity measurements it provides are critical for an E2E AEB model to make split-second, accurate decisions, and it can also ensure reliability at night when vision is compromised.
Vision-Language-Action (VLA) Benchmark. We will establish a full-chain, end-to-end benchmark for Vision-Language-Action (VLA) models. We will develop rich instruction sets, including scene description, object queries, occupancy status descriptions, and meta-actions, to ground language commands in the physical world and ultimately achieve end-to-end trajectory planning. Furthermore, this benchmark will encourage researchers to explore deeper multi-modal fusion strategies, particularly how to combine the precise motion information provided by the novel 4D radar to better understand action-related commands, thereby creating a complete evaluation loop from high-level semantic understanding to low-level trajectory planning.
Limitation
A limitation of the current version of OmniHD-Scenes is that the public road data is collected from a single geographical region (Changsha, China). While we have maximized the diversity of scenarios and road types within this extensive area, algorithms trained exclusively on this data may face challenges generalizing to regions with different road structures, traffic patterns, and driving behaviors. To efficiently address this, we plan to implement a targeted collection scheme guided by prior map knowledge to pinpoint distinct road infrastructures and regional traffic statistics to identify unique traffic cultures, thereby maximizing geographical diversity with optimized costs.
@@ Acknowledgement
Thanks to Intelligent Connected Vehicle inspection Center (Hunan) of CAERI Co., Ltd and Shanghai Geometrical Perception and Learning Co., Ltd.